OpenStudy (mrs.ambrose614):

WILL MEDAL Which of the following is equal to the square root of the cube root of 5? 5 to the power of 1 over 3 5 to the power of 1 over 6 5 to the power of 2 over 3 5 to the power of 3 over 2 CAN ANYONE HELP ME @518nad @zepdrix @mathmale @3mar @milo123 @help!!!!

OpenStudy (mrs.ambrose614):

Wouldn't it be b????

OpenStudy (mathmale):

Think\[\sqrt{(\sqrt[3]{5}})\]

OpenStudy (mathmale):

Your best bet would be to convert both into exponential form and then use rules of exponents.

OpenStudy (mrs.ambrose614):

How do I do that @mathmale @mathmath333 @milo123

OpenStudy (mathmale):

Problems such as these can be written in either exponential form or radical form. Don't you have a discussion of that in your textbook or online learning materials?

OpenStudy (mathmale):

Let a be a base. Then\[a ^{\frac{ b }{ c }}=\sqrt[c]{a^b}\]

OpenStudy (mathmale):

You must know this relationship.

OpenStudy (mathmale):

Therefore\[\sqrt[3]{5}=5^{\frac{ 1 }{ 3 }}\]

OpenStudy (mathmale):

That equation shows the radical form of "cube root of 5" on the left and the exponential form of "cube root of 5" on the right. Convert "square root of the cube root of 5" into exponential form. It will first appear like "square root of 5^(1/3), or

OpenStudy (mathmale):